# WebAccessVL: Making an Accessible Web via Violation-Conditioned VLM

### v1

@@ 5 Conclusion
Modify the following HTML code to comply with WCAG 2.2 accessibility
guidelines. Ensure the output is a properly formatted HTML file without
additional explanations or descriptions. Return only the modified HTML
code without extra text or commentary.
The expected output HTML has <zero/non-zero> violations.
@@
Inference details. During inference, we generate the output HTML code by using the same prompt as training. However, in order to conduct negative-guided sampling, we performed two forward passes, using $c_{=0}$ and $c_{>0}$, i.e., putting zero and non-zero in the condition.
@@
accessibility and WCAG 2.1 compliance. You are tasked with auditing and
@@
A5 Test Statistics of User Study
@@
A6 Training and Inference Time
@@
A7 Website Correction Process for WebAccessVL
@@
A8 Broader Impact
This work contributes to improving web accessibility by using vision-language models to automatically reduce WCAG2 violations in HTML. It has the potential to benefit users with disabilities and support developers by lowering the technical barrier to creating accessible websites. By relying on open-weight models, our approach also offers a privacy-conscious alternative to commercial services. However, automated fixes may optimize for rule-based compliance without ensuring real usability, and over-reliance on such systems could reduce the incentive for inclusive design practices. We view this method as a tool to support, not replace, human developers and encourage continued collaboration with accessibility experts and users.

### v2

@@ 5 Conclusion
The following HTML was checked for WCAG 2.2 accessibility compliance.
Violation information is unknown. / For each violation listed below,
locate the offending element and apply the fix directly in the HTML.
<Parsed Violation Report>
<cond> The expected output HTML has unknown violations.
Modify the HTML below. Return only the corrected HTML
without explanations.
@@
Inference details. During inference, we generate the output HTML code using the same prompt structure as training. To apply negative guidance sampling, we perform two forward passes at each token generation step: one conditioned on the violation report $\bm{c}$} and one with the unconditional prompt $\bm{c}$$_{\emptyset}$. The logits from both passes are combined using the guidance scale $\gamma$ to compute the final token probabilities.
For checker-in-the-loop inference, we iteratively refine the output for a maximum of $K=2$ passes. At each iteration, we: (1) generate corrected HTML ${\bm{y}}^{(i)}$ from the input conditioned on violations ${\bm{c}}^{(i-1)}$, (2) run the accessibility checker on ${\bm{y}}^{(i)}$ to obtain residual violations ${\bm{c}}^{(i)}$, and (3) if violations remain and $i<K$, use ${\bm{y}}^{(i)}$ and ${\bm{c}}^{(i)}$ as input for the next iteration. We track the HTML with the lowest violation count across all iterations as the final output. In practice, most samples converge after the first or second pass.
@@
accessibility and WCAG 2.2 compliance. You are tasked with auditing and
@@
A6 Test Statistics of User Study
@@
A7 Training and Inference Time
@@
A8 Website Correction Process for WebAccessVL
@@
A9 Broader Impact
This work contributes to improving web accessibility by using vision-language models to automatically reduce WCAG 2.2 violations in HTML. It has the potential to benefit users with disabilities and support developers by lowering the technical barrier to creating accessible websites. By relying on open-weight models, our approach also offers a privacy-conscious alternative to commercial services. However, automated fixes may optimize for rule-based compliance without ensuring real usability, and over-reliance on such systems could reduce the incentive for inclusive design practices. We view this method as a tool to support, not replace, human developers and encourage continued collaboration with accessibility experts and users.